Mixing equal parts of 5% (acidic)vinegar with 20% (acidic)vinegar will produce WHAT __% (acidic) vinegar?
Found 2 solutions by stanbon, richwmiller:
Answer by stanbon(75887)

Mixing equal parts of 5% (acidic)vinegar with 20% (acidic)vinegar will produce WHAT __% (acidic) vinegar?
-------
Note:: If you add x amt. to x amt. you end up with 2x amt.
--------------
Equation:
0.05x + 0.20x = 0.25x = (0.125)(2x)
----
Ans: 12.5 %
-------------------
Cheers,
Stan H.

Answer by richwmiller(17219)

5%*100+20%*100=x*200
1/20*100+1/5*100=x*200
5+20=x*200
25=200x
25/200=x
x=1/8=12.5%
another way
5%*100+20%*100=x*200
(.05+.20)*100=200x
.25=2x
x=.125=12.5%